You and your team are financial consultants who have been hired by a large, publicly-traded electronics firm, Brilliant Electronics (BI), a leader in its industry. The company is looking into manufacturing its new product, a machine using sophisticated state of the art technology developed by BI’s R&D team, overseas. This overseas project will last for five years. They’ve asked you to evaluate this project and to make a recommendation about whether or not the company should pursue it. BI’s management team needs your recommendation and the analysis used to arrive at it by no later than December 3, 2019.

The following market data on BI’s securities are current:

Debt: 210,000 6.4 percent coupon bonds outstanding, 25 years to maturity, selling or 108 percent of par; the bonds have $1000 par value each and make semi-annual payments

Common Stock: 8,300,000 shares outstanding, selling for $68 per share; beta=1.1

Preferred Stock: 450,000 shares of 4.5% preferred stock outstanding, selling or $81 per share

Market: 7 percent expected market risk premium; 3.5 percent risk-free rate

The company bought some land three years ago for $3.9 million in anticipation of using it as a toxic dump site for waste chemicals, but it built a piping system to safely discard the chemicals instead. The land was appraised last week for $4.4 million on an after-tax basis.   In five years, the after-tax value of the land will be $4.8 million, but the company expects to keep the land for a future project. The company wants to build its new manufacturing plant on this land; the plant will cost $37 million to build.

At the end of the project (the end of year 5), the plant can be scrapped for $5.1 million. The manufacturing plant will be depreciated using the straight-line method.

The company will incur $6,700,000 in annual fixed costs excluding depreciation. The plan is to manufacture 15,300 machines per year and sell them at $11,450 per machine; the variable production costs are $9,500 per machine. Selling price and costs are expected to remain unchanged over the life of the project.

BI uses PK Global (PKG) as its lead underwriter. PKG charges BI spreads of 8% on new common stock issues, 6% on new preferred stock issues, and 4% on new debt issues. PKG has included all direct and indirect issuance costs (along with its profit) in setting these spreads. BI’s tax rate is 35 percent. The project requires $1,300,000 in initial net working capital investment to get operational. Assume BI raises all equity for new projects externally (that is, BI does not use retained earnings).

The weighted average flotation cost is the sum of the weight of each source of funds in the capital structure of the company times the flotation costs, so:

fT = ($564.4/$827.65)(0.08) + ($36.45/$827.65)(0.06) + ($226.8/$827.65)(0.04) = 0.0682, or 6.82%

Thus the initial investment is increased by the amount of flotation costs:

                  (Amount raised)(1 – 0.0682) = $37,000,000   

                  Amount raised = $37,000,000/(1 – 0.0682) = $39,708,092

  1. Calculate the firm’s current cost of capital using the information provided.

  1. Calculate the project’s cost of capital (the appropriate discount rate to use to evaluate BI’s new project) assuming the capital structure will remain the same if the project is undertaken.

This project is somewhat riskier than a typical project for BI; therefore, management has asked you to use an adjustment factor of 12% to account for this increased riskiness (that is, to add 12% to the firm’s cost of capital) to estimate the project’s required rate of return.

(NOTE: Flotation costs do not have to be considered when calculating the required rate of return for each class of security – they are addressed in this problem by adjusting the cost of the initial investment to $39,708,092 from $37,000,000).
Answer #1

Calculation of WACC:

WACC = sum of weighted costs of capital

After-tax cost of debt: FV = 1,000; PV = -108%*FV = 108%*1,000 = 1,080; PMT (semi-annual coupon) = 6.4%*1,000/2 = 32; N = 25*2 = 50, solve for RATE.

Semi-annual YTM = 2.895% so annual YTM = 2*2.90% = 5.79%

After-tax cost of debt = 5.79%*(1-35%) = 3.76%

Cost of preferred stock = annual dividend/current price = (4.5%*1,000)/81 = 5.56%

Cost of common stock = risk-free rate + beta*market premium = 3.5% + (1.1*7%) = 11.20%

WACC does not have to be adjusted for flotation cost as flotation cost is adjusted in the initial investment required.

Formula n*p V/T Debt Preferred stock Common stock Total (T) Number of securities (n) 210000 450000 830000 Price per security

Current cost of capital for the firm = 8.91%

Cost of capital for the project = 20.91%

Note: Question does not mention if any further calculation of cash flows/NPV/IRR etc is required. If it is, please mention in comments and it will be done.
Answer #2

Hi I am currently working on this case for my finance, is there anything else you can share with me to assist me? My email is sstcanada@gmail.com ... It would be greatly appreciated!